Jim Murray and Phyllis Lowe received a total of $250,000 from a deceased relative's estate. They decided to put away $50,000 for their child and divide the remainder into 3/4 for Phyllis and 1/4 for Jim. How much will Phyllis and Jim receive?
Solve the unknown

Answer by checkley71(8403)

250,000-50,000=200,000 LEFT TO DIVIDE AMONG PHYLLIS & JIM.
3/4*200,000=150,000 FOR PHYLLIS.
200,000-150,000=50,000 FOR JIM.
PROOF
50,000+150,000+50,000=250,000
250,000=250,000